OpenStudy (anonymous):

Find the distance from the given point to the given line. Estimate your answer to the nearest 2 decimal places.

OpenStudy (anonymous):

did u need to find the shortest distance?

OpenStudy (anonymous):

yes, distance means shortest distance

OpenStudy (anonymous):

Can someone break this down for me?

OpenStudy (anonymous):

it will take 3 steps 1) find the equation of the line perpendicular to \(y=-2x-8\) through the point \((-2,6)\) 2) find the intersection of the two lines 3) find the distance between the point of intersection and the point \99-2,6)\) be happy to help with any of the steps

OpenStudy (anonymous):

Thank you! could you take me through them?

OpenStudy (anonymous):

sure for the first one, the line \(y=-2x-8\) has slope \(-2\) so the perpendicular line will have slope \(\frac{1}{2}\) the point slope formula gives \[y-6=\frac{1}{2}(x+2)\] \[y-6=\frac{1}{2}x+1\] \[y=\frac{1}{2}x+7\]

OpenStudy (anonymous):

for the second part, set \[\frac{1}{2}x+7=-2x-8\] and solve for \(x\) \[x+14=-4x-16\] \[5x=-30\] \[x=-6\] since \(x=-6\) we get \(y=-2\times -6-8=12-8=4\) so the point of intersection is \((-6,4)\)

OpenStudy (anonymous):

finally use the distance formula to find the distance between the points \((-6,4)\) and \((-2,6)\)

OpenStudy (anonymous):

\[\sqrt{(-6+2)^2+(6-4)^2}=\sqrt{4^2+2^2}=\sqrt{20}\]

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

k if you have a question about any step, ask

OpenStudy (anonymous):

can you do one more for me?

OpenStudy (anonymous):

why not

OpenStudy (anonymous):

ok thank you! and so for the first one it would be 20?

OpenStudy (anonymous):

no it is not 20, it is the square root of 20 i.e. \[\sqrt{20}\] or \[2\sqrt{5}\] they are the same number

OpenStudy (anonymous):

so how would I break it down?

OpenStudy (anonymous):

not sure what you mean the distance formula gives the answer as \(\sqrt{20}\) and since \(2=4\times 5\) we can write \[\sqrt{20}=\sqrt{4\times 5}=\sqrt{4}\sqrt{5}=2\sqrt{5}\]

OpenStudy (anonymous):

i mean since \(20=4\times 5\)

OpenStudy (anonymous):

if you want a decimal approximation, you need to use a calculator

OpenStudy (anonymous):

(-2, 6), y = -2x - 8

OpenStudy (anonymous):

this is what you wrote above right? find the distance between the point and the line

OpenStudy (anonymous):

yes I needed help

OpenStudy (anonymous):

i wrote all the steps above to find the distance. the final answer is \(\sqrt{20}\)

OpenStudy (anonymous):

no I mean the other lol

OpenStudy (anonymous):

other problem

OpenStudy (anonymous):

what is the other problem?

OpenStudy (anonymous):

(-2, 6), y = -2x - 8

OpenStudy (anonymous):

i think you wrote the same one down twice

OpenStudy (anonymous):

ohhh sorry this one

OpenStudy (anonymous):

(-6,4), y = -2x + 7

OpenStudy (anonymous):

same idea equation of the perpenicular line is \[y-4=\frac{1}{2}(x+6)\] \[y-4=\frac{1}{2}x+3\]\[y=\frac{1}{2}x+7\]

OpenStudy (anonymous):

set the two lines equal, get \[\frac{1}{2}x+7=-2x+7\] so \(x=0\) and therefore \(y=7\) for the point of intersection

OpenStudy (anonymous):

then find the distance between \((0,7)\) and \((-2,6)\) you get \[\sqrt{2^2+(7-6)^2}=\sqrt{5}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw, happy to help a tv star!